# Approximation properties of torsion classes

### v1

@@ Abstract.
We clarify some results of Bagaria and Magidor [MR3152715] about the relationship between large cardinals and torsion classes of abelian groups, and prove that (1) the Maximum Deconstructibility principle introduced in [Cox_MaxDecon] requires large cardinals; it sits, implication-wise, between Vopěnka’s Principle and the existence of an $\omega_{1}$-strongly compact cardinal. (2) While deconstructibility of a class of modules always implies the precovering property by [MR2822215], the concepts are (consistently) non-equivalent, even for classes of abelian groups closed under extensions, homomorphic images, and colimits.
@@ 1. Introduction
Classical homological algebra involves approximating an arbitrary module with modules from the well-behaved class $\mathcal{P}_{0}$ of projective modules. Relative homological algebra ([MR178036]) attempts to replace $\mathcal{P}_{0}$ by some other well-behaved class $\mathcal{C}$, leading to new invariants that can be used to create new module and ring theoretic invariants. A crucial requirement for this to work nicely is that $\mathcal{C}$ be a precovering class in the sense of Enochs and Auslander (see Section 2 for definitions).
It can be difficult to show that a class is precovering, but Saorín and Šťovíček proved that
[…]
where deconstructibility is a concept that arose from the solution of the Flat Cover Conjecture ([MR1798574], [MR1832549]) and is closely related to Quillen’s Small Object Argument from homotopy theory. A natural question arose:
@@
Suppose $\mathcal{C}$ is a class of modules that is closed under transfinite extensions. Does the converse of (italic-$\star$1) hold?
After clarifying some results of Bagaria and Magidor from [MR3152715], we use their results to provide a negative answer to Question 1, at least in the absence of large cardinals:
@@
is always a covering class that is closed under transfinite extensions, homomorphic images, and colimits. We prove:
@@
Maximum Deconstructibility appeared very powerful, since [Cox_MaxDecon] showed that it implied deconstructibility of many classes in Gorenstein Homological Algebra that (so far) are not known to be deconstructible in ZFC alone. But it was unclear whether Maximum Deconstructibility had any large cardinal strength at all. We again use the (clarified) results of Bagaria and Magidor to show that it does:
@@
Section 2 has preliminaries. Section 3 proves the approximation properties of torsion classes that are provable in ZFC alone. Section 4 clarifies some results of Bagaria-Magidor [MR3152715]. Section 5 proves the main theorems mentioned above, and Section 6 includes some questions.
@@ 3. Approximation properties and torsion classes
    $\mathcal{K}$ is $(\kappa,\lambda)$-cofinal if $\mathcal{K}\cap[M]^{<\kappa}$ is $\subseteq$-cofinal in $[M]^{<\lambda}$ for every module $M\in\mathcal{K}$ (i.e., whenever every $N\in[M]^{<\lambda}$ is contained in some $L\in\mathcal{K}\cap[M]^{<\kappa}$).
@@
    $\mathcal{K}$ is bounded by $\kappa$ whenever $M=\sum\left(\mathcal{K}\cap[M]^{<\kappa}\right)$ for all $M\in\mathcal{K}$ (i.e., if every $x\in M$ is contained in some ${<}\kappa$-presented submodule of $M$ that lies in $\mathcal{K}$). [^2] This property was introduced by Gardner in [MR0647168] and later investigated by Dugas in [MR0799895].
@@
Clearly any $(\kappa,\lambda)$-cofinal class is bounded by $\kappa$. In the forthcoming Lemma 11 we will argue that for certain classes $\mathcal{K}$, the concepts are equivalent.
@@
4. Clarification of some results of Bagaria and Magidor
In Bagaria-Magidor [MR3152715, §5] a torsion class ${}^{\perp_{0}}\mathcal{X}$ is called “$\kappa$-generated” whenever every $A\in{}^{\perp_{0}}\mathcal{X}$ is a direct sum of subgroups in ${}^{\perp_{0}}\mathcal{X}$, each of cardinality ${<}\kappa$; this is what is usually called $\kappa$-decomposable (see Definition 4). However, their use of the word “direct” on page 1867 of [MR3152715] appears to be a misprint, since the arguments there never use (or conclude) directness of the relevant sums. All of the arguments in [MR3152715, §5], and the citation they provide for the concept (Dugas [MR0799895]), appear to use the weaker property that every group in the class is merely a sum of subgroups in the class of cardinality ${<}\kappa$. The latter condition is what we called bounded by $\kappa$ in Section 2 (this terminology was used by Gardner [MR0647168] and Dugas [MR0799895]).
The proof of [MR3152715, Theorem 5.1] shows that if $\kappa$ is a $\omega_{1}$-strongly compact cardinal and $X$ is a countable abelian group, then ${}^{\perp_{0}}X$ is bounded by $\kappa$. In fact, they prove the stronger property of ${}^{\perp_{0}}X$ being $(\kappa,\omega_{1})$-cofinal. [^3] Recall that this stands for the following property: For each $G\in{}^{\perp_{0}}\mathbb{Z}$ every countable subgroup $H\leq G$ is included in a member of ${}^{\perp_{0}}\mathbb{Z}\cap[G]^{<\kappa}$ (see Definition 4). The proof however does not show that ${}^{\perp_{0}}X$ is $\kappa$-decomposable, and in fact that is impossible, because of the following theorem (due to the third author):
@@ 4. Clarification of some results of Bagaria and Magidor
The proof of their Theorem 5.1 really showed:
Theorem 8 (clarification of [MR3152715, Theorem 5.1]).
If $\kappa$ is a $\delta$-strongly compact cardinal and $X$ is an abelian group of cardinality less than $\delta$, then ${}^{\perp_{0}}X$ is bounded in $\kappa$.
In the other direction, Theorem 5.3 of [MR3152715] asserted that if ${}^{\perp_{0}}\mathbb{Z}$ is “$\kappa$-generated” (using their definition, with “direct sum” instead of just “sum”), then $\kappa$ is $\omega_{1}$-strongly compact. Fortunately, their argument never used the “direct” part of their definition, since by Theorem 7, that would have been an inconsistent assumption. Their proofs really showed:
Theorem 9 (clarification of [MR3152715, Theorem 5.3]).
Corollary 10 (clarification of [MR3152715, Corollary 5.4]).
@@ 5. Maximum Deconstructibility, and deconstructibility of torsion classes, require large cardinals
Through this section $\kappa$ will denote a regular uncountable cardinal. Our main goal is to clarify, for a given class of modules $\mathcal{K}$, the relationship between $\mathcal{K}$ being $\kappa$-deconstructible and $\mathcal{K}$ being bounded in $\kappa$. As noted in §2, any $(\kappa,\kappa)$-cofinal class $\mathcal{K}$ is bounded in $\kappa$. In fact the former property seems to be strictly stronger than the latter, and they both follow from $\kappa$-deconstructibility. Lemma 11 provides some important scenarios where these concepts to be equivalent. Combining this lemma with the (clarified) results of Bagaria and Magidor from Section 4 allows us to prove the main results mentioned in the introduction, namely:
@@
Now suppose $\mathcal{K}$ is closed under transfinite extensions and quotients. We prove (b) $\Rightarrow$ (a). Without losing any generality we may assume that (the universe of) every module $M\in\mathcal{K}$ is a cardinal. Since $\mathcal{K}$ is closed both under transfinite extensions and quotients, then by [Cox_MaxDecon, Theorem 6.3] in order to prove $\kappa$-deconstructibility of $\mathcal{K}$ it suffices to show that if $M\in\mathcal{K}$, then $\mathcal{K}\cap\wp^{*}_{\kappa}(M)$ is stationary in $\wp^{*}_{\kappa}(M)$, where $\wp^{*}_{\kappa}(M)$ denotes the ${<}\kappa$-sized submodules of $M$ whose intersection with $\kappa$ is transitive.
The set $\mathcal{K}\cap\wp^{*}_{\kappa}(M)$ being stationary in $\wp^{*}_{\kappa}(M)$ is equivalent to the following statement: For each algebra $\mathfrak{A}=(M,\dots)$ on $M$ in a countable signature, there is $X\prec\mathfrak{A}$ in $\mathcal{K}\cap\wp^{*}_{\kappa}(M)$. We prove this next.
Using the assumption that $\mathcal{K}$ is $\subseteq$-cofinal in $[M]^{<\kappa}$, and the downward Löwenheim-Skolem Theorem, we construct an $\subseteq$-increasing $\omega$-chain
$X_{0}\subseteq C_{0}\subseteq X_{1}\subseteq C_{1}\subseteq\dots$
@@
  • $X_{n}\prec\mathfrak{A}$, $|X_{n}|<\kappa$, and $X_{n}$ has transitive intersection with $\kappa$.
  • $C_{n}\in\mathcal{K}$ and $|C_{n}|<\kappa$ (but might have non-transitive intersection with $\kappa$).
Let $X:=\bigcup_{n<\omega}X_{n}$. Then $X\prec\mathfrak{A}$, $|X|<\kappa$, and $X\cap\kappa$ is transitive. Also, note that $X=\bigcup_{n<\omega}C_{n}$ too. Since $\mathcal{K}$ is closed under quotients, each $C_{n+1}/C_{n}$ is in $\mathcal{K}$. So since $\mathcal{K}$ is closed under transfinite extensions, and each $C_{n+1}/C_{n}$ is in $\mathcal{K}$, we conclude that $X\in\mathcal{K}$.
@@
We now focus on abelian groups. Recall that by Lemma 6 any torsion class ${}^{\perp_{0}}\mathcal{X}$ is closed under homomorphic images, transfinite extensions, and colimits. So by Lemma 11, a class of the form ${}^{\perp_{0}}\mathcal{X}$ is deconstructible if and only if it is bounded in some cardinal. Furthermore, as noted in the introduction, although an $\omega_{1}$-strongly compact cardinal might be singular, all cardinals above it are also $\omega_{1}$-strongly compact. In particular (by considering its successor if necessary), there exists an $\omega_{1}$-strongly compact cardinal if and only if there exists a regular $\omega_{1}$-strongly compact cardinal. Then Lemma 6, Lemma 11, and the (clarified) Corollary 10 of Bagaria and Magidor immediately imply Theorem 2 from the introduction, which asserted that ${}^{\perp_{0}}\mathbb{Z}$ is deconstructible if and only if there is an $\omega_{1}$-strongly compact cardinal. In fact, making use of the (clarified) Theorem 8 of Bagaria and Magidor, we have:
@@ 6. Questions
Question 15.
Suppose that $\kappa\geq\omega_{2}$ is a regular cardinal such that ${}^{\perp_{0}}X$ is deconstructible for all abelian group of size ${<}\kappa$. Must $\kappa$ be strongly compact?

### v2

@@ Abstract.
We strengthen a result of Bagaria and Magidor [MR3152715] about the relationship between large cardinals and torsion classes of abelian groups, and prove that (1) the Maximum Deconstructibility principle introduced in [Cox_MaxDecon] requires large cardinals; it sits, implication-wise, between Vopěnka’s Principle and the existence of an $\omega_{1}$-strongly compact cardinal. (2) While deconstructibility of a class of modules always implies the precovering property by [MR2822215], the concepts are (consistently) non-equivalent, even for classes of abelian groups closed under extensions, homomorphic images, and colimits.
2020 Mathematics Subject Classification:
16E30, 16S90, 03E75, 16D90, 18G25, 16B70
@@ 1. Introduction
Approximation theory of modules provides tools for studying general modules by choosing special classes of modules and finding approximations (precovers and preenvelopes), and minimal approximations (covers and envelopes), of the general modules by modules in the chosen classes.
A model case is the special class $\mathcal{P}_{0}$ of projective modules that yields projective precovers, and projective resolutions of modules. Similarly, the class $\mathcal{F}_{0}$ of flat modules yields flat covers, and minimal flat resolutions of modules, [MR1832549]. Dually, the class $\mathcal{I}_{0}$ of injective modules yields injective envelopes, and minimal injective coresolutions of modules.
In the case when minimal approximations exist, they provide invariants of modules, the model examples being Bass’ invariants and dual Bass’ invariants of modules over commutative noetherian rings [MR1753146].
Approximation theory makes it possible to extend classic homological algebra to the setting of relative homological algebra [EilMoo]. Applications of the extended setting are numerous: in the theory of finitely generated modules over artin algebras [AR], tilting theory of commutative noetherian rings [MR2985554], and Gorenstein homological algebra [MR3839274], to name a few.
Classic structure theory of modules is based on decompositions into (possibly infinite) direct sums of indecomposable modules. This is a strong tool in the study of modules of finite length and injective modules over commutative noetherian rings [AF]. However, decomposability fails for many other classes of interest. In contrast, the weaker notion of deconstructibility (namely, decomposition into a transfinite extension of modules from a given set of modules) is almost omnipresent.
The key point is that deconstructible classes of modules always provide precovers, that is, each deconstructible class is precovering [MR2822215]. The opposite question is more subtle:
@@
Suppose $\mathcal{C}$ is a class of modules that is precovering and closed under transfinite extensions. Is $\mathcal{C}$ deconstructible?
In this paper we use some results of Bagaria and Magidor from [MR3152715] to provide a negative answer to Question 1, at least in the absence of large cardinals:
@@
is always a covering class that is closed under transfinite extensions, homomorphic images, and colimits. In this respect we prove:
@@
Maximum Deconstructibility appeared very powerful, since [Cox_MaxDecon] showed that it implied deconstructibility of many classes in Gorenstein Homological Algebra that (so far) are not known to be deconstructible in ZFC alone. But it was unclear whether Maximum Deconstructibility had any large cardinal strength at all. We again use the results of Bagaria and Magidor to show that it does:
@@
Section 2 has preliminaries. Section 3 proves the approximation properties of torsion classes that are provable in ZFC alone. Section 4 discusses some results of Bagaria-Magidor [MR3152715]. Section 5 proves the main theorems mentioned above, and Section 6 includes some questions.
@@ 3. Approximation properties and torsion classes
    $\mathcal{K}$ is $(\kappa,\lambda)$-cofinal if $\mathcal{K}\cap[M]^{<\kappa}$ is $\subseteq$-cofinal in $[M]^{<\lambda}$ for every module $M\in\mathcal{K}$ (i.e., whenever every $N\in[M]^{<\lambda}$ is contained in some ${<}\kappa$-generated submodule of $M$ that lies in $\mathcal{K}$).
@@
    $\mathcal{K}$ is bounded by $\kappa$ whenever $M=\sum\left(\mathcal{K}\cap[M]^{<\kappa}\right)$ for all $M\in\mathcal{K}$ (i.e., if every $x\in M$ is contained in some ${<}\kappa$-generated submodule of $M$ that lies in $\mathcal{K}$). [^2] This property was introduced by Gardner in [MR0647168] and later investigated by Dugas in [MR0799895].
@@
Clearly any $(\kappa,\lambda)$-cofinal class is bounded by $\kappa$. In the forthcoming Lemma 11 we will argue that for certain classes $\mathcal{K}$, “bounded by $\kappa$” and ``$(\kappa,\kappa)$-cofinal” are equivalent concepts.
@@
4. On some results of Bagaria and Magidor
In Bagaria-Magidor [MR3152715, §5] a torsion class ${}^{\perp_{0}}\mathcal{X}$ is called “$\kappa$-generated” whenever every $A\in{}^{\perp_{0}}\mathcal{X}$ is a direct sum of subgroups in ${}^{\perp_{0}}\mathcal{X}$, each of cardinality ${<}\kappa$; this is what is usually called $\kappa$-decomposable (see Definition 4). The use of the word “direct” on page 1867 of [MR3152715] appears to be a misprint, [^3] The anonymous referee has suggested that the authors of [MR3152715] might have intended to say directed sum instead of direct sum. since the arguments there never use (or conclude) directness of the relevant sums. The arguments in [MR3152715, §5], and the citation they provide for the concept (Dugas [MR0799895]), appear to use the weaker property that every group in the class is a sum of subgroups in the class of cardinality ${<}\kappa$. The latter condition is what we called bounded by $\kappa$ in Section 2 (this terminology was used by Gardner [MR0647168] and Dugas [MR0799895]). So we shall use the bounded by $\kappa$ terminology to state the Bagaria-Magidor results. They proved:
Theorem 7 ([MR3152715, Theorem 5.1]).
If $\kappa$ is a $\delta$-strongly compact cardinal and $X$ is an abelian group of cardinality less than $\delta$, then ${}^{\perp_{0}}X$ is bounded by $\kappa$. In fact, they prove the stronger property of ${}^{\perp_{0}}X$ being $(\kappa,\omega_{1})$-cofinal. [^4] Recall that this stands for the following property: For each $G\in{}^{\perp_{0}}\mathbb{Z}$ every countable subgroup $H\leq G$ is included in a member of ${}^{\perp_{0}}\mathbb{Z}\cap[G]^{<\kappa}$ (see Definition 4).
Theorem 8 ([MR3152715, Theorem 5.3]).
Corollary 9 ([MR3152715, Corollary 5.4]).
The following ZFC theorem (due to the third author) shows that, while ${}^{\perp_{0}}\mathbb{Z}$ can be bounded (by the Bagaria-Magidor results), it can never be decomposable:
@@ 5. Maximum Deconstructibility, and deconstructibility of torsion classes, require large cardinals
Through this section $\kappa$ will denote a regular uncountable cardinal. Our main goal is to clarify, for a given class of modules $\mathcal{K}$, the relationship between $\mathcal{K}$ being $\kappa$-deconstructible and $\mathcal{K}$ being bounded in $\kappa$. As noted in §2, any $(\kappa,\kappa)$-cofinal class $\mathcal{K}$ is bounded in $\kappa$. In fact the former property seems to be strictly stronger than the latter, and they both follow from $\kappa$-deconstructibility. Lemma 11 provides some important scenarios where these concepts are equivalent. Combining this lemma with the results of Bagaria and Magidor from Section 4 allows us to prove the main results mentioned in the introduction, namely:
@@
Now suppose $\mathcal{K}$ is closed under transfinite extensions and quotients. We prove (b) $\Rightarrow$ (a). Without losing any generality we may assume that (the universe of) every module $M\in\mathcal{K}$ is a cardinal. Since $\mathcal{K}$ is closed both under transfinite extensions and quotients, and $|R|<\kappa$, then by [Cox_MaxDecon, Theorem 6.3] in order to prove $\kappa$-deconstructibility of $\mathcal{K}$ it suffices to show that if $M\in\mathcal{K}$, then $\mathcal{K}\cap\wp^{*}_{\kappa}(M)$ is stationary in $\wp^{*}_{\kappa}(M)$, where $\wp^{*}_{\kappa}(M)$ denotes the ${<}\kappa$-sized submodules of $M$ whose intersection with $\kappa$ is transitive (i.e., those $<\kappa$-sized submodules $N$ of $M$ such that $N\cap\kappa\in\kappa$).
The set $\mathcal{K}\cap\wp^{*}_{\kappa}(M)$ being stationary in $\wp^{*}_{\kappa}(M)$ is equivalent to the following statement ([MattHandbook]): for each function $F$ from finite subsets of $M$ into $M$, there is an $X\in\mathcal{K}\cap\wp^{*}_{\kappa}(M)$ such that $X$ is closed under the function $F$. We prove this next.
Fix an arbitrary function $F$ from finite subsets of $M$ into $M$; notice that since $F$ is finitary, for any infinite $D\subset M$, the closure of $D$ under the function $F$ has the same cardinality as $D$. Using the assumption that $\mathcal{K}$ is $\subseteq$-cofinal in $[M]^{<\kappa}$, fix some $C_{0}\in\mathcal{K}\cap[M]^{<\kappa}$. Recursively build a $\subseteq$-increasing $\omega$-chain
$C_{0}\subseteq X_{0}\subseteq C_{1}\subseteq X_{1}\subseteq C_{2}\subseteq X_{2}\subseteq\dots$
@@
  • $X_{n}$ is the closure of $C_{n}\cup\text{sup}(C_{n}\cap\kappa)$ under the function $F$. This closure has cardinality $<\kappa$, because $|C_{n}|<\kappa$ and hence (by regularity of $\kappa$) $|C_{n}\cup\text{sup}(C_{n}\cap\kappa)|<\kappa$.
  • (for $n\geq 1$) $C_{n}\in\mathcal{K}$, $C_{n}\supseteq X_{n-1}$, and $|C_{n}|<\kappa$. This is possible by the assumption that $\mathcal{K}$ is $\subseteq$-cofinal in $[M]^{<\kappa}$ (but note that $C_{n}\cap\kappa$ might fail to be transitive).
Let $X:=\bigcup_{n<\omega}X_{n}$. Then $X$ is closed under $F$, $|X|<\kappa$, and $X\cap\kappa=\text{sup}_{n<\omega}\text{sup}(C_{n}\cap\kappa)$; in particular, $X\cap\kappa$ is transitive. Also, note that $X=\bigcup_{n<\omega}C_{n}$. Since $\mathcal{K}$ is closed under quotients, each $C_{n+1}/C_{n}$ is in $\mathcal{K}$. So since $\mathcal{K}$ is closed under transfinite extensions, and each $C_{n+1}/C_{n}$ is in $\mathcal{K}$, we conclude that $X\in\mathcal{K}$.
@@
We now focus on abelian groups. Recall that by Lemma 6 any torsion class ${}^{\perp_{0}}\mathcal{X}$ is closed under homomorphic images, transfinite extensions, and colimits. So by Lemma 11, a class of the form ${}^{\perp_{0}}\mathcal{X}$ is deconstructible if and only if it is bounded in some cardinal. Furthermore, as noted in the introduction, although an $\omega_{1}$-strongly compact cardinal might be singular, all cardinals above it are also $\omega_{1}$-strongly compact. In particular (by considering its successor if necessary), there exists an $\omega_{1}$-strongly compact cardinal if and only if there exists a regular $\omega_{1}$-strongly compact cardinal. Then Lemma 6, Lemma 11, and the Corollary 9 of Bagaria and Magidor immediately imply Theorem 2 from the introduction, which asserted that ${}^{\perp_{0}}\mathbb{Z}$ is deconstructible if and only if there is an $\omega_{1}$-strongly compact cardinal. In fact, making use of the Theorem 7 of Bagaria and Magidor, we have:
@@
Remark 15.
The following statement has recently been proven in [PPT, Remark 4 and Corollary 3.4]: Assume Vopěnka’s Principle holds. Then for any ring $R$, each class of $R$-modules that is closed under finite direct sums, extensions, and direct limits is deconstructible.
@@ 6. Questions
Question 16.
Suppose that $\kappa\geq\omega_{2}$ is a regular cardinal such that ${}^{\perp_{0}}X$ is deconstructible for all abelian groups of size ${<}\kappa$. Must $\kappa$ be strongly compact?